OpenStudy (anonymous):

simplify the complex fraction (x/x+2)/(1/x)+(1/x+2)

OpenStudy (anonymous):

x/(x+2) /(1/x) + (1/(x+2) is that correct

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

3x +2 +1/x

OpenStudy (anonymous):

not one of the answers

OpenStudy (anonymous):

x/(x+2) * x/1 = x^2 /(x+2)

OpenStudy (anonymous):

choices are x^2/2 or x^2/x+2 or x^2/2x+2 or x^2+2x+2/2x+2

OpenStudy (anonymous):

x^2 /(x+2) + 1/(x+2) =x^2 +1 /(x+2)

OpenStudy (anonymous):

that is an option. what did you use for the common denom?

OpenStudy (anonymous):

no that isn't one of the choice

OpenStudy (anonymous):

I don't see a response. i saw you were writing.

OpenStudy (anonymous):

um ok i was doing it over to see where i went wrong but im not seeing a mistake when you divide (x/x+2) by (1/x) you should get x^2 / x+2

OpenStudy (anonymous):

ok that is one of my choices, I thought you had x^2 +1 /x+2

OpenStudy (anonymous):

yea i did cus when you add 1/ x+2 you would get the answer i had before

OpenStudy (anonymous):

ok got it thanks